OpenStudy (anonymous):

find the indefinite integral ∫(sec 2x+ tan 2x) dx

OpenStudy (anonymous):

please help. cal 1 class

OpenStudy (math&ing001):

\[\int\limits_{}^{}(\sec 2x + \tan 2x)=\int\limits_{}^{}\sec 2x dx +\int\limits_{}^{} \tan 2x dx\] As for the first see math2.org/math/integrals/more/sec.htm For the second \[\int\limits_{}^{} \tan 2x dx=-\frac{ 1 }{ 2 } \int\limits_{}^{}\frac{ -2\sin 2x }{ \cos 2x }dx=-\frac{ 1 }{ 2 }Ln(\cos(2x))\]